# Multilingual Synopses of Movie Narratives: A Dataset for Story Understanding

### v1

@@ Abstract
Story video-text alignment, a core task in computational story understanding, aims to align video clips with corresponding sentences in their descriptions. However, progress on the task has been held back by the scarcity of manually annotated video-text correspondence and the heavy concentration on English narrations of Hollywood movies. To address these issues, in this paper, we construct a large-scale multilingual video story dataset named Multilingual Synopses of Movie Narratives (M-SyMoN), containing 13,166 movie summary videos from 7 languages, as well as manual annotation of fine-grained video-text correspondences for 101.5 hours of video. Training on the human annotated data from M-SyMoN outperforms the SOTA methods by 15.7 and 16.2 percentage points on Clip Accuracy and Sentence IoU scores, respectively, demonstrating the effectiveness of the annotations. As benchmarks for future research, we create 6 baseline approaches with different multilingual training strategies, compare their performance in both intra-lingual and cross-lingual setups, exemplifying the challenges of multilingual video-text alignment.
Multilingual Synopses of Movie Narratives: A Dataset for Story Understanding
@@ 1 Introduction
Computational story understanding aims to empower AI systems with the ability to delve into the intricacies of diverse stories, unlocking their deep semantics such as character motivations and intentions Emelin et al. (2021); Rashkin et al. (2018), event structures Chambers and Jurafsky (2008); Du et al. (2021), and social relationships among story characters Elson et al. (2010); Chaturvedi et al. (2016); Kim and Klinger (2019). In recent years, computation story understanding has garnered significant research interest Dong et al. (2023); Sang et al. (2022); Andrus et al. (2022); Xu et al. (2022); Han et al. (2023) and many story understanding tasks Wu and Krahenbuhl (2021); Choi et al. (2021) have emerged.
@@
Figure 1: An example alignment between a video clip sequence and a sentence sequence from M-SyMoN. One text chuck may correspond to several video clips, while some video clips may not match any textual description.
@@
We further investigate the multi-lingual characteristics of the dataset, and make the following observations. First, in the intra-lingual setup, compared to translating all languages to English for training and inference, additional language-specific fine-tuning on weakly supervised data brings an average improvement of 5.4 percentage points across 7 languages. Adding a small portion of the manual annotations further boosts performance. Second, in the cross-lingual setup, for source and target languages that are linguistically similar (e.g., Spanish to Portuguese or French), the transfer performance is generally good; for source and target languages that are different (e.g., Chinese to English or Hindi), the transfer performance is quite limited. Third, an out-of-domain evaluation on the YMS dataset shows that training on the weakly supervised data from M-SyMoN outperforms the state of the art methods by 12.3 and 13.2 percentage points on Clip Accuracy and Sentence IoU scores. Moreover, adding manually annotated video-text alignment data further improves the performance by 2.4 and 3.0 percentage points, indicating the utility of our annotated alignment data.
@@ 2.2 Movie Story Datasets
Although there are many story understanding dataset, multilingual video story datasets are scarce. To our knowledge, the only dataset is Movie101v2 Yue et al. (2024), which contains 46K Chinese video-caption pairs from 203 movies, where the Chinese text are translated to English using GPT-3.5-turbo. The dataset also lacks human-annotated video-text correspondence.
In comparison, M-SyMoN is large-scale and multilingual, containing movie summary videos in 7 languages and 101.5 hours of video with exact video-text correspondences.
@@ 2.3 Multilingual Story Understanding
Methods for multilingual story understanding mainly fall into two categories: (1) translating training/test data to English Ponti et al. (2020) or prompts to the target language Lin et al. (2022), and (2) directly fine-tuning on non-English text using a multilingual Pretrained Language Model (PLM) Tikhonov et al. (2021) or a PLM for the target language Lai et al. (2022). In this paper, we benchmark M-SyMoN on multilingual story video-text alignment with both types of methods.
@@ 3 Task Formulation
Number of Annotated Videos 98 31 75 84 57 63 72
@@ 4.1 Data Source and Statistics
For data collection, we first identify YouTube channels with movie recap videos in the target languages. Keywords such as “movie summary in <language>” and “movie recap in <language>” are used to search for these channels. We then download all videos and their accompanying subtitles from the identified channels. Videos without subtitles and that are not movie summaries are discarded.
This yields a total of 13,166 videos or 2,136 video hours in 7 languages, including English, Chinese, Spanish, French, Portuguese, Hindi, and Russian. We list the statistics in Table 1. M-SyMoN contains summaries for 5,960 movies and TV shows, of which 1,515 have more than one summary.
@@ 4.5 Dataset Analysis
As most movie metadata are not available from the movie recap videos, we use ChatGPT (gpt-3.5-turbo-0125) to acquire the meta information. Specifically, we input the video title, the video-level description featured on YouTube, and first 5 sentences of the movie narration into ChatGPT and ask ChatGPT to output the movie title and metadata.
@@ 5.1 Base Model
For every clip-text pair, we randomly sample $K$ hard negatives from the same video. With the sampled negative text features and video features, we fine-tune the encoders by minimizing the contrastive InfoNCELoss (Oord et al., 2018) below:
@@
During inference, we first obtain the representation of each clip and sentence, and then calculate the cosine similarity between every clip-sentence pair. Finally, we find the global clip-sentence alignment from the similarity scores of each pair by resorting to a sequence alignment algorithm named Double Drop Dynamic Time Warp (Drop-DTW) Dvornik et al. (2021), detailed in Appendix D.1.
@@
5.2 Multilingual Fine-Tuning Methods
@@
  • Individual training (CCLM-individual). Using weakly supervised training data, we fine-tune the base model on each language and obtain 7 language-specific models.
  • Translational training (CCLM-translate). We first translate all the weakly supervised training data and the test data in M-SyMoN to English with off-the-shelf translation model NLLB-200 Costa-jussà et al. (2022). We then fine-tune the base model on the translated data to obtain a unified model.
  • Two-stage training (CCLM-two-stage). On top of the trained CCLM-translate model, we further fine-tune on each language separately to obtain 7 language-specific models.
To show the usefulness of our manually annotated video-text alignment dataset, we further fine-tune two aforementioned methods on the human annotated training set in Section 4.4:
  • CCLM-translate-supervision. We translate the non-English languages in the human annotated data to English and fine-tune the CCLM-translate model on the translated data.
  • CCLM-two-stage-supervision. We fine-tune the CCLM-two-stage model on the human annotated data of each language separately.
@@ 6.1 Experimental Setup
For CCLM-translate and CCLM-multilingual, we initialize their parameters from the pre-trained CCLM-base model Zeng et al. (2023) and fine-tune them for 20 epochs with an initial learning rate of $4\times 10^{-5}$ and cosine learning rate decay. For CCLM-two-stage, we initialize its parameters from the CCLM-translate model and fine-tune it for 20 epochs with an initial learning rate of $4\times 10^{-6}$ and cosine learning rate decay. Random augmentation and weight decay of 0.2 are applied.
@@ 6.2 Evaluation Metrics
Following Dogan et al. (2018), we use two evaluation metrics. Clip accuracy is defined as the temporal proportion of correctly aligned video segments. Sentence IoU is defined as the intersection-over-union between the aligned video durations and the groundtruth durations. Due to space limitations, we mainly report the F1 score, i.e. the harmonic mean between Clip Accuracy and Sentence IoU in this section and defer detailed results to Appendix D.2.
@@ 6.3 Intra-Lingual Results
First, we find that training the base model on all languages together performs significantly worse than training on each language separately. Specifically, CCLM-individual outperforms CCLM-multilingual by 4.3-17.3 percentage points. This suggests that language-specific features are important for multilingual video-text alignment. Second, the CCLM-translate baseline is on par with CCLM-individual and significantly outperforms CCLM-multilingual. Moreover, the two-stage baseline (i.e., CCLM-two-stage) leverages the benefits of pre-training on a large multilingual dataset and subsequent fine-tuning on a specific language dataset, which outperforms CCLM-individual and CCLM-translate by 3.1 and 5.4 percentage points on average. Finally, after finetuning with our manually annotated data, it yields consistent improvements on top of both CCLM-translate and CCLM-two-stage models across the 7 languages, and CCLM-two-stage-supervision achieves the best average performance among all the compared methods. These observations demonstrate the usefulness of the manually annotated video-text alignment data.
@@
Figure 5: Case Study. The original movie descriptions are in Chinese. The CCLM-translate baseline is trained and evaluated on the text translated to English with NLLB-200 Costa-jussà et al. (2022). The CCLM-two-stage baseline is evaluated on the Chinese text, and we show the human-translated English text here for clear illustration. Green boxes mark correct alignments and red boxes mark incorrect alignments.
@@ 6.4 Cross-Lingual Transfer Results
We employ the CCLM-two-stage model to evaluate the cross-lingual performance of every source-target language pair. The results are reported in Figure 4.
@@
First, the highest value of each column appears on the diagonal. This is intuitive because the model is trained and evaluated on the same language. Second, linguistically similar languages transfer well to each other. For example, related languages like Spanish and Portuguese, generally obtain good cross-lingual performance. This likely stems from shared vocabulary, sub-word tokens, and grammatical structures. Furthermore, Chinese does not transfer well to any other language, possibly due to differences in linguistic construction and movie distributions (See §4.5) . Lastly, we can observe that it is difficult to transfer from any language to Hindi, possibly because the pre-training data of CCLM does not contain Hindi, which hurts representation learning.
6.5 Ablation: Transfer to YMS
@@
In this section, we extend our analysis to show that models trained on M-SyMoN can generalize to other video-text alignment benchmarks. Specifically, we evaluate baseline models trained on M-SyMoN on the out-of-domain English benchmark dataset YMS Dogan et al. (2018) and compare the results with the state-of-the-art performances.
@@
As shown in Table 4, the large scale M-SyMoN dataset significantly improves the performance of YMS. First, we can observe that fine-tuning the base model on the human-annotated training set of YMS (i.e., CCLM-individual (YMS)) is already on par with the state-of-the-art performance, demonstrating the effectiveness of our proposed base model. Second, compared to CCLM-individual (YMS), only training on the weakly supervised portion of M-SyMoN improves performance by 13.7 and 12.7 percentage points on Clip Accuracy and Sentence IoU, respectively. This suggests the value of the large-scale weakly-supervised data in M-SyMoN. Moreover, additional fine-tuning on the human-annotated data from M-SyMoN, i.e., CCLM-two-stage-supervision (English), further boosts the performance by 2.4 and 3.0 percentage points on Clip Accuracy and Sentence IoU, respectively. These observations highlight the significance and generalizability of our human-annotated video-text correspondence dataset, indicating the potential of M-SyMoN for advancing research in video story understanding.
@@
Table 4: Out-of-domain results on the YMS dataset. CCLM-individual (YMS) is the base model fine-tuned on the human-annotated training set of YMS.
6.6 Case Study
In Figure 5, we provide a qualitative example highlighting the challenges of M-SyMoN.
First, the movie descriptions are intended to advance the story plot and do not describe the video content exactly, making alignment difficult. Second, the sentences and clips in a movie are very similar causing difficulty in alignment. For example, Figure 5 shows a war film with numerous scenes of car bombings. Without understanding the story and merely judging by the visual and textual content, the description “Vehicles set on fire during an attack” can be aligned with any of these scenes. Finally, the translation model may incur errors. For example, the last sentence in the second line of Figure 5 contains translation errors and is incomprehensible. CCLM-translate mistakenly aligns this sentence to a scene of the gunfire because “attack” appears in the sentence, while CCLM-two-stage, with access to the correct Chinese caption, makes the correct prediction.
@@ 7 Conclusion
In this paper, we introduced M-SyMoN, a large-scale multilingual video story dataset. It contains 13,166 movie summary videos from 7 languages, featuring 480 videos with human-annotated video-text correspondence. We established several multilingual multimodal baselines to benchmark M-SyMoN. Experimental results demonstrate the value of M-SyMoN for video story understanding. Its size, multilingual nature, and rich alignment annotations make M-SyMoN a valuable contribution.

### v2

# Multilingual Synopses of Movie Narratives: A Dataset for Vision-Language Story Understanding

@@ Abstract
Story video-text alignment, a core task in computational story understanding, aims to align video clips with corresponding sentences in their descriptions. However, progress on the task has been held back by the scarcity of manually annotated video-text correspondence and the heavy concentration on English narrations of Hollywood movies. To address these issues, in this paper, we construct a large-scale multilingual video story dataset named Multilingual Synopses of Movie Narratives (M-SyMoN), containing 13,166 movie summary videos from 7 languages, as well as manual annotation of fine-grained video-text correspondences for 101.5 hours of video. Training on the human annotated data from M-SyMoN outperforms the SOTA methods by 15.7 and 16.2 percentage points on Clip Accuracy and Sentence IoU scores, respectively, demonstrating the effectiveness of the annotations. As benchmarks for future research, we create 6 baseline approaches with different multilingual training strategies, compare their performance in both intra-lingual and cross-lingual setups, exemplifying the challenges of multilingual video-text alignment. The dataset is released at: https://github.com/insundaycathy/M-SyMoN
Multilingual Synopses of Movie Narratives: A Dataset for Vision-Language Story Understanding
@@ 1 Introduction
Computational story understanding aims to empower AI systems with the ability to delve into the intricacies of diverse stories, unlocking their deep semantics such as character motivations and intentions Emelin et al. (2021); Rashkin et al. (2018), event structures Chambers and Jurafsky (2008); Du et al. (2021), and social relationships among story characters Elson et al. (2010); Chaturvedi et al. (2016); Kim and Klinger (2019). In recent years, computation story understanding has garnered significant research interest Dong et al. (2023); Sang et al. (2022); Andrus et al. (2022); Xu et al. (2022); Han et al. (2023) and many story understanding tasks Wu and Krahenbuhl (2021); Choi et al. (2021); Gu et al. (2023) have emerged.
@@
Figure 1: An example alignment between a video clip sequence and a sentence sequence from M-SyMoN. One text chuck may correspond to several video clips, while some video clips may not match any textual description. The example is from https://youtu.be/n5v9hzSYxPQ, a summary of the movie 407 Dark Flight (2012)
@@
We further investigate the multi-lingual characteristics of the dataset, and make the following observations. First, in the intra-lingual setup, compared to translating all languages to English for training and inference, additional language-specific finetuning on weakly supervised data brings an average improvement of 5.4 percentage points across 7 languages. Adding a small portion of the manual annotations further boosts performance. Second, in the cross-lingual setup, for source and target languages that are linguistically similar (e.g., Spanish to Portuguese or French), the transfer performance is generally good; for source and target languages that are different (e.g., Chinese to English or Hindi), the transfer performance is quite limited. Third, an out-of-domain evaluation on the YMS dataset shows that training on the weakly supervised data from M-SyMoN outperforms the state of the art methods by 12.3 and 13.2 percentage points on Clip Accuracy and Sentence IoU scores. Moreover, adding manually annotated video-text alignment data further improves the performance by 2.4 and 3.0 percentage points, indicating the utility of our annotated alignment data.
@@ 2.2 Movie Story Datasets
Although there are many story understanding dataset, multilingual video story datasets are scarce. To our knowledge, the only dataset is Movie101v2 Yue et al. (2024), which contains 46K Chinese video-caption pairs from 203 movies, where the Chinese text are translated to English using GPT-3.5-turbo. The dataset also lacks human-annotated video-text correspondence. In comparison, M-SyMoN is large-scale and multilingual, containing movie summary videos in 7 languages and 101.5 hours of video with exact video-text correspondences.
@@ 2.3 Multilingual Story Understanding
Methods for multilingual story understanding mainly fall into two categories: (1) translating training/test data to English Ponti et al. (2020) or prompts to the target language Lin et al. (2022), and (2) directly finetune on non-English text using a multilingual Pretrained Language Model (PLM) Tikhonov et al. (2021) or a PLM for the target language Lai et al. (2022). In this paper, we benchmark M-SyMoN on multilingual story video-text alignment with both types of methods.
@@ 3 Task Formulation
In the task formulation, we make the simplifying assumption that multiple video clips can be aligned with one sentence, but a clip cannot be aligned with more than one sentence. The reason is that a video clip is typically very short (around 2.4 seconds) and clips outnumber sentences, so it is rare that one clip aligns with more than one sentence.
@@
Number of Annotated Videos 98 30 75 84 57 63 72
@@ 4.1 Data Source and Statistics
For data collection, we first identify YouTube channels with movie recap videos in the target languages. Keywords such as “movie summary in <language>” and “movie recap in <language>” are used to search for the channels. We then download all videos and their accompanying subtitles from the identified channels. Videos without subtitles and that are not movie summaries are discarded.
This yields 13,166 videos or 2,136 video hours in 7 languages, including English, Chinese, Spanish, French, Portuguese, Hindi, and Russian. We list the statistics in Table 1. M-SyMoN contains summaries for 5,960 movies and TV shows, of which 1,515 have more than one summary.
@@ 4.5 Dataset Analysis
As most movie metadata are not available from the movie recap videos, we use ChatGPT (gpt-3.5-turbo-0125) to acquire the information. Specifically, we input the video title, the video-level description featured on YouTube, and first 5 sentences of the movie narration into ChatGPT and ask ChatGPT to output the movie title and metadata.
@@ 5.1 Base Model
For every clip-text pair, we randomly sample $K$ hard negatives from the same video. With the sampled negative text features and video features, we finetune the encoders by minimizing the contrastive InfoNCELoss (Oord et al., 2018) below:
@@
During inference, we first obtain the representation of each clip and sentence, and then calculate the cosine similarity between every clip-sentence pair. Finally, we find the global clip-sentence alignment from the similarity scores of each pair by resorting to a sequence alignment algorithm named Double Drop Dynamic Time Warp (Drop-DTW) Dvornik et al. (2021), detailed in Appendix E.1.
@@
5.2 Multilingual Finetuning Methods
@@
  • Individual training (CCLM-individual). Using weakly supervised training data, we finetune the base model on each language and obtain 7 language-specific models.
  • Translational training (CCLM-translate). We first translate all the weakly supervised training data and the test data in M-SyMoN to English with off-the-shelf translation model NLLB-200 Costa-jussà et al. (2022). We then finetune the base model on the translated data to obtain a unified model.
  • Two-stage training (CCLM-two-stage). On top of the trained CCLM-translate model, we further finetune on each language separately to obtain 7 language-specific models.
To show the usefulness of our manually annotated video-text alignment dataset, we further finetune two aforementioned methods on the human annotated training set in Section 4.4:
  • CCLM-translate-supervision. We translate the non-English languages in the human annotated data to English and finetune the CCLM-translate model on the translated data.
  • CCLM-two-stage-supervision. We finetune the CCLM-two-stage model on the human annotated data of each language separately.
@@ 6.1 Experimental Setup
For CCLM-translate and CCLM-multilingual, we initialize their parameters from the pre-trained CCLM-base model Zeng et al. (2023) and finetune them for 20 epochs with an initial learning rate of $4\times 10^{-5}$ and cosine learning rate decay. For CCLM-two-stage, we initialize its parameters from the CCLM-translate model and finetune it for 20 epochs with an initial learning rate of $4\times 10^{-6}$ and cosine learning rate decay. Random augmentation and weight decay of 0.2 are applied.
@@ 6.2 Evaluation Metrics
Following Dogan et al. (2018), we use two evaluation metrics. Clip accuracy is defined as the temporal proportion of correctly aligned video segments. Sentence IoU (Jaccard, 1908) is defined as the intersection-over-union between the aligned video durations and the groundtruth durations. Due to space limitations, we mainly report the F1 score, i.e. the harmonic mean between Clip Accuracy and Sentence IoU in this section and defer detailed results to Appendix E.2.
@@ 6.3 Intra-Lingual Results
First, we find that training the base model on all languages together performs significantly worse than training on each language separately. Specifically, CCLM-individual outperforms CCLM-multilingual by 4.3-17.3 percentage points. This suggests that language-specific features are important for multilingual video-text alignment. Second, the CCLM-translate baseline is on par with CCLM-individual and significantly outperforms CCLM-multilingual. Moreover, the two-stage baseline (i.e., CCLM-two-stage) leverages the benefits of pre-training on a large multilingual dataset and subsequent finetuning on a specific language dataset, which outperforms CCLM-individual and CCLM-translate by 3.1 and 5.4 percentage points on average. Finally, after finetuning with our manually annotated data, it yields consistent improvements on top of both CCLM-translate and CCLM-two-stage models across the 7 languages, and CCLM-two-stage-supervision achieves the best average performance among all the compared methods. These observations demonstrate the usefulness of the manually annotated video-text alignment data.
@@ 6.4 Cross-Lingual Transfer Results
We employ the CCLM-two-stage model to evaluate the cross-lingual performance of every source-target language pair, see Figure 4 for results.
@@
First, the highest value of each column appears on the diagonal. This is intuitive because the model is trained and evaluated on the same language. Second, linguistically similar languages transfer well to each other. For example, related languages like Spanish and Portuguese, generally obtain good cross-lingual performance. This likely stems from shared vocabulary, sub-word tokens, and grammatical structures. Furthermore, Chinese does not transfer well to any other language, possibly due to differences in linguistic construction and movie distributions (See §4.5) . Lastly, we observe it is difficult to transfer from any language to Hindi, possibly because the CCLM pre-training data does not contain Hindi, which hurts representation learning.
English Chinese Spanish French Portuguese Hindi Russian Average
Without Timestamps of Text
CCLM-two-stage-supervision 27.7 38.9 19.8 21.0 19.5 12.9 17.5 22.5
With Timestamps of Text
Transcript baseline 34.4 42.6 19.1 26.2 40.6 20.1 27.3 30.1
CCLM-2S-supervision-time 34.6 46.2 28.5 27.7 39.8 25.4 26.4 32.7
Table 4: Intra-lingual video-text alignment results with transcript temporal information.
6.5 Effects of Timestamps of Narration Text
All previous models do not utilize the timestamps of the narration texts as input features or during inference. As such timestamps are only available after a video summary is made, models relying on timestamps as input features cannot handle general alignment tasks, such as aligning plot summaries with movies.
In this section, we explore the effectiveness of these timestamps during inference. Specifically, we use the CCLM-two-stage-supervised model to calculate video-text similarity. During DTW, we constrain video clips to match only sentences within a 1-second window of their timestamps. This model is named CCLM-2S-supervised-time. We also create a transcript baseline that relies solely on the timestamps; we simply align video clips to the sentence whose timestamp falls in their duration.
Table 4 summarizes the results. First, the transcript baseline achieves an average F1 score of only 30.1, indicating significant discrepancies between the temporal transcripts and our annotations. This demonstrates the noisy nature of YouTube transcripts and underscores the importance of precise annotations for accurate evaluation. Second, CCLM-2S-supervised-time outperforms CCLM-two-stage-supervised by 10.2 percentage points, showing that transcript temporal information improves alignment. Finally, CCLM-2S-supervised-time surpasses the transcript baseline by 2.6 percentage points, demonstrating that our model learns video-text correspondences beyond the transcripts.
6.6 Transfer to YMS
@@
In this section, we extend our analysis to investigate if models trained on M-SyMoN can generalize to other video-text alignment benchmarks, such as the out-of-domain English benchmark dataset YMS Dogan et al. (2018).
@@
As shown in Table 5, the large scale M-SyMoN dataset significantly improves the performance of YMS. First, we can observe that finetuning the base model on the human-annotated training set of YMS (i.e., CCLM-individual (YMS)) is already on par with the state-of-the-art performance, demonstrating the effectiveness of our proposed base model. Second, compared to CCLM-individual (YMS), only training on the weakly supervised portion of M-SyMoN improves performance by 13.7 and 12.7 percentage points on Clip Accuracy and Sentence IoU, respectively. This suggests the value of the large-scale weakly-supervised data in M-SyMoN. Moreover, additional finetuning on the human-annotated data from M-SyMoN, i.e., CCLM-two-stage-supervision (English), further boosts the performance by 2.4 and 3.0 percentage points on Clip Accuracy and Sentence IoU, respectively. These observations highlight the significance and generalizability of our human-annotated video-text correspondence dataset, indicating the potential of M-SyMoN for advancing research in video story understanding.
@@
Table 5: Out-of-domain results on the YMS dataset. CCLM-individual (YMS) is the base model finetuned on the human-annotated training set of YMS.
6.7 Discussion
The experiments in this paper demonstrate the value of M-SyMoN as a multilingual and multimodal story understanding dataset. First, M-SyMoN is the first dataset that offers large-scale human annotations for multilingual story video-text alignment, and such annotations deliver pratical benefits. As shown in §6.5, the timestamps in the YouTube transcripts are noisy, and human annotations are necessary for accurate benchmarking. Additionally, the human-annotated training set provides valuable supervision signals. In §6.3, we achieve a notable 3.2% relative improvement from human annotation data that amounts to only 2.2% of the training data. Moreover, finetuning on supervised data from M-SyMoN lead significant improvement on the out-of-domain YMS dataset in §6.6. Finally, M-SyMoN provides fertile ground for linguistic and cultural studies of story elements. For example, cross-lingual transfer experiments (§6.4) lends support to existing linguistic theories on positive and negative transfer in language learning Eronen et al. (2023); Odlin (1989) as linguistically similar languages transfer well to each other. The analysis of movie themes for different languages (Figure 3) demonstrates the cultural relevance of M-SyMoN.
@@ 7 Conclusion
In this paper, we introduced M-SyMoN, a large-scale multilingual video story dataset. It contains 13,166 movie summary videos from 7 languages, featuring 480 videos with human-annotated video-text correspondence. We established several multilingual multimodal baselines to benchmark M-SyMoN. Experimental results show the value of M-SyMoN for video story understanding. Its size, multilingual nature, and rich alignment annotations make M-SyMoN a valuable contribution.
Acknowledgments
We gratefully acknowledge the support by the Nanyang Associate Professorship and the National Research Foundation Fellowship (NRF-NRFF13-2021-0006), Singapore. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not reflect the views of the funding agencies.